(I'm struggling with this problem because, I can calculate the interest but I don't know how to figure out the deposit)

Suppose you want to save in order to purchase a new boat. Take the APR to be 8.4%. If you deposit $200 each month, how much will you have toward the purchase of a boat after three years? (Round your answer to the nearest cent.)
Answer #1

Future Value=Monthly amount/(APR/12)*((1+APR/12)^(12*t)-1)=200/(8.4%/12)*((1+8.4%/12)^(12*3)-1)=8156.2006634